OpenStudy (anonymous):

Area. The perimeter of a rectangle is 197 inches. If the rectangle is x inches wide, write a function for the area A(x). I'm a little bit confused any help?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (jhannybean):

Perimeter of a rectangle can be found by using the formula \[\large P=2l+2w\] where l=length and w=width. They give you the width. width = w = x and P= 197. So our new equation for perimeter after inputting all the variables would be \[\large 197=2l+2x\] we can solve for \(l\) by subtracting 2x from both sides and dividing by 2. We will get \[\large l=\frac{-2x+197}{2}= -x +\cfrac{197}{2}\] Now we can substitute \(l\) into the equation for Area of a rectangle, which is : \(\large A = l \cdot w\) Substituting \(l\) back into our equation, we get \[\large A = \left(-x+\cfrac{197}{2}\right)(x)\]\[\large A=-x^2 +\cfrac{197}{2}x\]

OpenStudy (anonymous):

thanks so much. i really do appreciate your help and explanations...i had 2421ft^2 which is the area....but i don't think they were asking that

OpenStudy (jhannybean):

Yeah just a general equation